# PAL: Probing Audio Encoders via LLMs - Audio Information Transfer into LLMs

### v2

@@ Abstract
Integration of audio perception into large language models (LLMs) is an emerging research area for enabling machine listening applications, yet efficient transfer of rich audio semantics from audio encoders to LLMs remains underexplored. The most widely used integration paradigm projects the audio encoder output tokens into the LLM input space (e.g., via an MLP or a Q-Former), then prepends or inserts them to the text tokens. We refer to this generic scheme as Prepend to the LLM’s input token space (PLITS) integration. We propose an efficient alternative, Lightweight Audio LLM Integration (LAL). LAL introduces audio representations solely via the attention mechanism within different layers of the LLM, bypassing its feedforward module. LAL encodes rich audio semantics at an appropriate level of abstraction for integration into different blocks of LLMs. Our design significantly reduces computational overhead compared to existing integration approaches. Observing with Whisper that the speech encoder benefits from PLITS integration, we propose an audio encoder aware approach for efficiently Probing Audio encoders via LLM (PAL), which employs PLITS integration for Whisper and LAL for general audio encoders. Under an identical training curriculum, LAL consistently maintains performance or outperforms existing integration approaches across multiple base LLMs and tasks. For general audio tasks, LAL improvement is up to 30% over a strong PLITS baseline while reducing memory usage by up to 64.1% and increasing throughput by up to 247.5%. Furthermore, for general audio-music-speech LLM, PAL, performs on par with a fully PLITS integration-based system but with substantially improved computational and memory efficiency. Project page:https://ta012.github.io/PAL/
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2024; Liu et al., 2024a) have emerged as the foundational technology for natural language interaction with machines, demonstrating remarkable conversational fluency. Despite this success, their perceptual capabilities remain limited primarily to text, restricting their ability to understand the physical world. This limitation has inspired significant research into multi-modal LLMs (MLLMs), which expand traditional LLMs by integrating additional sensory modalities such as vision (Vision LLMs) (Liu et al., 2023; Templeton et al., 2024; Wang et al., 2024), audio (Large Audio Language Models (LALMs) or simply audio-LLMs) (Deshmukh et al., 2023; Gong et al., 2024; Tang et al., 2024; Ghosh et al., 2024; 2025a), and other inputs (Brohan et al., 2023; Thawkar et al., 2023) to foster more natural, intuitive, and effective human-machine interfaces.
Recent advances in audio representation learning have produced powerful encoders trained with self-supervised objectives  Huang et al. (2022); Alex et al. (2025) and multimodal supervised objectives (CLAP Elizalde et al. (2023); Wu et al. (2023). We argue that the primary function of an audio LLM is to query the audio encoder’s representations via natural language and return the information the user wants. This requires an integration mechanism that reliably transfers key audio information, including event cues and temporal dynamics, from the encoder into the LLM embedding and concept space; once received, the LLM can apply its reasoning and generation to answer queries from concise facts to nuanced creative text. Our work focuses on  optimizing this information transfer.
An audio-LLM typically consists of (i) a large language model (LLM), (ii) one or more audio encoders, and (iii) an integration strategy that connects encoder outputs to the LLM. Regarding the selection of audio encoders, prior audio-LLM work often performs additional pretraining of a chosen encoder to extend its capabilities (Goel et al., 2025; Chu et al., 2024b; Gong et al., 2024). For example, the Whisper encoder (Radford et al., 2023), originally trained for speech-to-text transcription, is further adapted for general audio event understanding in (Goel et al., 2025); the AST encoder (Gong et al., 2022a; 2021) is retrained for language alignment in (Gong et al., 2024; Ghosh et al., 2024). In short, these approaches select a single encoder and add new abilities through further pretraining. In contrast, our approach leverages multiple  off the shelf audio encoders that were trained on different domains like general audio(sound), speech etc. and with diverse training mechanisms, including self-supervised learning with training objectives to capture fine-grained information (Huang et al., 2022; Ahmed et al., 2024; Chen et al., 2024; Alex et al., 2025), language aligned contrastive training (Elizalde et al., 2023; Wu et al., 2023; Ghosh et al., 2025b), and transcription based next text token prediction (Radford et al., 2023). Reusing such pretrained encoders avoids redundant pretraining, promotes reuse across communities working on self supervised learning, speech recognition, and CLAP driven multimodal alignment, and improves efficiency in audio LLM development.
@@
When it comes to the integration of audio encoders with the LLM, two architectural paradigms dominate today. The first transforms the outputs of an audio encoder or encoders into the LLM input space (e.g., via an MLP, a QFormer (Li et al., 2023), etc.), then prepend or insert these audio tokens to the text tokens and propagates the entire sequence through all LLM layers as if decoding jointly over audio and text. Please note that the common theme in this family is how audio tokens are passed to the LLM: they are prepended to the text tokens. We refer to this generic scheme Prepend to the LLM’s input token space (PLITS) integration, a term we have introduced to group many state of the art methods in this family of audio LLMs such as  Wu et al. (2025b); Xu et al. (2025b); Chu et al. (2024b); Goel et al. (2025); Chu et al. (2023a); Ghosh et al. (2024); Tang et al. (2024); Gong et al. (2024); Deshmukh et al. (2023). The second paradigm, Flamingo style architectures (Alayrac et al., 2022; Kong et al., 2024), instead insert cross attention and feedforward (FFN) blocks between successive LLM layers; at each insertion, text tokens attend to a set of latent audio tokens, pass through the block FFN, and only then proceed to the next LLM layer. While this design improves attention efficiency relative to PLITS concatenation, the interleaved cross attention plus FFN stacks increase sequential depth and per layer compute, which can slow the forward pass.
@@
In contrast, we introduce LAL, a lightweight integration that injects audio tokens into the LLM’s attention blocks as keys and values only (without forming audio queries) and bypasses the LLM FFNs for audio tokens. This reduces the attention complexity from $\mathcal{O}\!\big((N_{a}{+}N_{t})^{2}\big)$ to $\mathcal{O}\!\big((N_{a}{+}N_{t})N_{t}\big)$, where $N_{a}$ and $N_{t}$ denote the numbers of audio and text tokens, respectively. Since typically $N_{a}\gg N_{t}$, this yields substantial efficiency gains. By avoiding both quadratic attention over audio tokens and their passage through the LLM FFNs, LAL achieves significant reductions in memory usage and computation. Unlike parameter efficient training methods such as LoRA, this is a core architectural modification, so the efficiency benefits are realized not only during training but also at inference time.
@@
Figure 1: Comparison of compute efficiency between LAL (ours) and PLITS, state of the art audio-LLM integration(our baseline). Training was performed with batch size 8 on an NVIDIA A100 using bfloat16, and inference with batch size 12 on an NVIDIA A100 using float16. All benchmarks were executed sequentially on the same node to eliminate load-related discrepancies.
PLITS and Flamingo integration techniques represent complementary strategies for extracting information from audio encoders. LAL provides a compute and memory efficient mechanism by limiting how audio tokens interact with the LLM, while other encoders may benefit from the richer decoding that occurs within the LLM under PLITS style integration. In particular, encoders trained with language contrastive or self supervised objectives such as CLAP and SSLAM are better served by LAL integration, whereas Whisper, which is pretrained with an autoregressive speech that is spoken language transcription and next token prediction objective, gains from the additional decoding capacity of PLITS style integration. Motivated by this observation, we propose a hybrid LAL plus PLITS framework called PAL for building general purpose audio, music, and speech LLMs, enabling encoder aware fusion that balances efficiency with performance. This design achieves strong results while substantially reducing computational and memory requirements compared to using PLITS style integration alone.
@@
To validate these architectural choices, we conduct a systematic empirical study under a standardized training curriculum and dataset setup, ensuring fair comparisons across models. Our experiments explore the trade-off between performance and efficiency, highlighting how encoder-aware fusion facilitates effective information transfer from audio encoders to LLMs with minimal parameter overhead. This analysis provides actionable insights into the design of scalable and efficient audio LLMs that leverage diverse pretrained audio encoders.
Our main contributions are as follows:
@@
1. We introduce LAL, a lightweight integration strategy for audio-LLMs that incorporates audio tokens solely as keys and values in the LLM’s attention sub-modules and skips FFNs, thereby reducing computation and memory cost while retaining performance comparable to PLITS integration.
2. We design an encoder-aware integrated LLM PAL that selectively applies LAL or PLITS integration depending on the audio encoder, enabling general-purpose audio, speech, and music LLMs that balance efficiency with performance.
3. We conduct fair and rigorous architectural comparisons under a standardized training curriculum and dataset setup, providing actionable insights into the efficiency–performance trade-offs of audio-LLM design.
@@ 2 Literature review
Audio LLM architectures: When integrating audio encoders with an LLM, two paradigms dominate. In PLITS, encoder features are mapped to the LLM token space with a small projector such as an MLP or a Q Former, the resulting audio tokens are typically prepended to the text tokens, and the joint sequence is processed by all LLM layers (Wu et al., 2025b; Xu et al., 2025b; Chu et al., 2024b; Goel et al., 2025; Chu et al., 2023a; Ghosh et al., 2024; Tang et al., 2024; Gong et al., 2024; Deshmukh et al., 2023). In contrast, the Flamingo style architecture inserts cross attention and feed forward adapters between successive LLM layers so that text tokens attend to latent audio tokens at selected depths (Alayrac et al., 2022; Kong et al., 2024). This makes audio to text interaction explicit and gated, but adds sequential depth, per layer compute, and parameters.
Audio-LLM Datasets: Beyond architecture, recent works curate high-quality instruction tuning datasets, both open source and proprietary (Goel et al., 2025; Ghosh et al., 2024; Chu et al., 2024b; Xu et al., 2025a) and build audio reasoning benchmarks (Sakshi et al., 2024; Deshmukh et al., 2025a; b). Training PLITS or Flamingo style models on these resources improves instruction following and scales reasoning across diverse audio tasks, with most gains attributable to these data innovations.
@@ 3 Methodology
This section outlines our approach to integrating audio with language models. We begin by formalizing PLITS, the SOTA audio-LLM integration, as our reference baseline. We then introduce LAL, a lightweight alternative that injects audio through attention only, and we analyze its compute and memory profile. Next, we present the experimental setup and results for classification, captioning, and reasoning, including scaling and a frozen FFN variant. Finally, we connect these findings to PAL, an encoder aware hybrid that selects between PLITS and LAL on a per encoder basis in order to support speech understanding without sacrificing efficiency on general audio.
3.1 Terminologies
We summarize the key terminologies used throughout the paper.
We use PLITS, the SOTA integration as our baseline, LAL as our proposed method, and PAL as a hybrid of the two. Please note that we use  LAL and  PAL to denote the integration approach and the corresponding audio-LLM. We employ SSLAM and CLAP, using an efficient Q-former-based connector that combines information from both inspired by  Tong et al. (2024) without increase in token count, referred to as LFST. When LFST is not used, the audio encoder defaults to SSLAM; when LFST is used, it represents a combination of SSLAM and CLAP. See Appendix E.1 for further details on LFST. Unless otherwise specified, we use Llama 3.2 1B Instruct (Grattafiori et al., 2024) as the base LLM. For evaluating larger models, we report Llama 3.2 3B Instruct  (Grattafiori et al., 2024) , and to assess transfer across model families, we also include Qwen2.5 1.5B Instruct (Team, 2024) .
@@ 3.2 Baseline Audio LLM: Prepend to the LLM’s input token space (PLITS)
To establish a fair comparison point for our integration methods, we construct a baseline audio LLM that follows the widely adopted SOTA integration Prepend to the LLM’s input token space (PLITS) paradigm. In this design, the audio encoder outputs are first mapped into the LLM input embedding space using a Q-Former–style connector. The resulting audio tokens are then prepended to the text tokens, and the concatenated sequence is propagated through all LLM layers so that decoding proceeds jointly over audio and text (see Fig. 2(A)). The central characteristic of this paradigm is how audio tokens are provided to the LLM: they are prepended to the text tokens. This integration strategy is used by most audio LLMs, including several state of the art systems Wu et al. (2025b); Xu et al. (2025b); Chu et al. (2024b); Goel et al. (2025); Chu et al. (2023a); Ghosh et al. (2024); Tang et al. (2024); Gong et al. (2024); Deshmukh et al. (2023).
@@ 3.3 LAL: Lightweight Audio-LLM Integration
A standard LLM layer contains an attention submodule followed by a feed-forward(FFN) submodule. Because attention mediates inter token interaction, it is the necessary conduit for audio to influence text, and we posit it is also sufficient for text tokens to gather information from audio. Guided by this principle, we introduce LAL (Lightweight Audio LLM integration). A shared Q Former produces a sequence of audio tokens as in our baseline, and at each layer a MLP projects these tokens into that layer’s input space. Audio information is then injected into the attention block only through Keys and Values while Queries remain text only, so audio modulates the attention context of text tokens without passing through the feed forward network.
@@
after which $\tilde{H}_{l}^{t}$ proceeds through the FFN with the usual residual connections. In this way, audio cues shape the attention context seen by text tokens, aligning audio-evoked features with their linguistic counterparts and enabling effective cross–modal information transfer.
Compute and memory efficiency.
LAL improves efficiency over PLITS and Flamingo style along three axes, and the benefits grow with longer audio sequences. We observe up to 64.1% lower memory usage and up to 247.5% higher training throughput (samples/sec). See Figure 1 for detailed training and inference metrics.
A. Attention complexity.
@@
LAL: audio tokens do not enter the feedforward sublayer and only serve as keys and values for text queries, which reduces per layer floating point operations and the activations stored for backpropagation.
@@
Scaling with audio length. Non text modalities in multimodal LLMs often yield far more tokens, and audio is no exception. As $N_{a}$ grows due to longer clips or denser tokenization, PLITS incurs a cost of $(N_{a}+N_{t})^{2}$, so the $N_{a}^{2}$ term dominates. In contrast, LAL scales as $(N_{a}+N_{t})N_{t}$, which is linear in $N_{a}$. Thus the compute and memory gap widens with longer or more finely segmented audio. The feedforward savings in LAL also increase with $N_{a}$ because a larger share of tokens bypass the most expensive part of each block.
Not PEFT or LoRA. (Hu et al., 2022)
@@
LAL is a core architectural change, not a parameter efficient fine tuning(PEFT) method. Techniques such as LoRA modify how weights are adapted during training while keeping the forward compute pattern essentially the same at inference. LAL changes attention and feedforward routing, so its compute and memory savings hold at inference as well as during training.
LAL Integration with Frozen LLM FFN. We also verify that LAL integration remains effective when the LLM’s FFN blocks are frozen, with no significant loss in performance (refer to Appendix E). This finding has important implications for reducing training cost, improving parameter efficiency, and preserving the pretrained knowledge of the LLM while enabling multimodal alignment. For clarity and consistency, however, our main experiments focus on the standard setting with trainable FFN blocks, and discussion of the frozen-FFN variant is limited to Appendix E.
Leveraging parametric versus contextual knowledge: Here we posit how LAL efficiently utilizes two types of knowledge inherent in pre-trained LLMs: (1) parametric knowledge, primarily embedded within the FFN layers as a result of extensive language pre-training, and (2) contextual knowledge, which is dynamically incorporated through attention mechanisms. The empirical success of LAL(refer to Table 1, Table  2) shows that audio input, as contextual information, can effectively induce required concept activations in text token representations via attention-based modulation, without needing direct FFN processing of audio representations. Consequently, audio information indirectly accesses the LLM’s parametric knowledge: the audio context ”piggybacks” on text tokens, as attention mechanisms reconfigure these representations, which then engage relevant concept-related pathways during FFN processing. Such a strategy offers gains in architectural efficiency and provides deeper insights into multimodal information integration.
Figure 2: Illustration of integration techniques: (A) SOTA integration PLITS (prepend to the LLM’s input token space), which prepends audio tokens to text tokens and propagates the full sequence through all LLM layers (our baseline); (B) our proposed lightweight integration LAL, which introduces audio representations only through the attention mechanism (see Equations 2, 3, and 4) while bypassing the feedforward modules; (C) the hybrid PAL, an encoder aware integration that combines LAL and PLITS by selecting the method for each encoder.
Empirical Evaluation of LAL We train and evaluate LAL on general audio tasks, including classification, captioning, and reasoning, across multiple base LLMs following the protocol in Section 4.1. To clearly separate contributions, we present two sets of results. First, in Table 1 (classification and captioning) and Table 2 (reasoning), we report a controlled comparison between LAL and PLITS, showing that LAL achieves comparable or better accuracy while being more efficient in speed and memory. Second, in Table 3 (classification and captioning) and Table 4 (reasoning), we compare LAL with prior works. Note that training data scale and model size vary significantly across prior approaches; our model operates on the lower end of both dimensions. These results should therefore be interpreted as evidence that LAL remains competitive despite using fewer resources.
@@
Table 1: Performance evaluation of the proposed efficient integration method LAL and SOTA integration PLITS across different base LLMs. Evaluation follows the protocol of Gong et al. (2024). AC: Audio caps, CL:Clotho AS2M: AudioSet 2M indicates CIDEr and indicates SPICE. Other metrics: accuracy (ESC-50, VocalSound), Mi-F1 (DCASE), and mAP (FSD, AudioSet). Complete evaluation methodology explained in Section 4.1 and dataset details in Appendix D
LLM PLITS LAL LFST Classification Captioning
@@
Table 2: GPT-4 evaluation of LAL and PLITS on the CompA-R benchmark (Ghosh et al., 2024). A text only GPT-4 judge scores the model outputs; see Ghosh et al. (2024) for the detailed prompt.

### v3

@@ Abstract
Integration of audio perception into large language models (LLMs) is an emerging research area for enabling machine-listening applications, yet efficient transfer of rich audio semantics from audio encoders to LLMs remains underexplored. The most widely used integration paradigm projects audio-encoder output tokens into the LLM input space (e.g., via an MLP or a Q-Former) and then prepends or inserts them into the text token sequence. We refer to this generic scheme as Prepend to the LLM’s input token space (PLITS) integration. We propose an efficient alternative, Lightweight Audio LLM Integration (LAL). LAL injects audio representations solely through the attention mechanism at selected LLM layers, bypassing the feed-forward module. It encodes rich audio semantics at an appropriate level of abstraction for integration into different transformer blocks, substantially reducing computational overhead compared to existing approaches. We further introduce PAL, a hybrid integration approach for efficiently Probing Audio encoders via LLM. PAL applies PLITS only to a compact set of summary tokens while integrating the full audio token sequence via LAL. Under an identical training curriculum, LAL consistently matches or outperforms existing integration approaches across multiple base LLMs and tasks, with improvements of up to 30% over a strong PLITS baseline, while reducing memory usage by about 60% and increasing throughput by about 190%. Moreover, PAL matches or exceeds PLITS performance while offering substantially better computational and memory efficiency.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2024; Liu et al., 2024a) have emerged as the foundational technology for natural language interaction with machines, demonstrating remarkable conversational fluency. Despite this success, their perceptual capabilities remain limited primarily to text, restricting their ability to understand the physical world. This limitation has inspired significant research into multi-modal LLMs (MLLMs), which expand traditional LLMs by integrating additional sensory modalities such as vision (Vision LLMs) (Liu et al., 2023; Templeton et al., 2024; Wang et al., 2024), audio (Large Audio Language Models (LALMs) or simply audio-LLMs) (Deshmukh et al., 2023; Gong et al., 2024; Tang et al., 2024; Ghosh et al., 2024, 2025a), and other inputs (Brohan et al., 2023; Thawkar et al., 2023) to foster more natural, intuitive, and effective human-machine interfaces.
@@
An audio LLM typically comprises three components: (i) a large language model (LLM), (ii) audio encoder that convert raw audio to audio features or tokens, and (iii) a mechanism that integrates encoder outputs into the LLM. When it comes to the integration of audio encoders with the LLM, two architectural paradigms dominate today. The first transforms the outputs of an audio encoder or encoders into the LLM input space (e.g., via an MLP, a QFormer (Li et al., 2023), etc.), then prepend or insert these audio tokens to the text tokens and propagates the entire sequence through all LLM layers as if decoding jointly over audio and text. Please note that the common theme in this family is how audio tokens are passed to the LLM: they are prepended to the text tokens. We refer to this generic scheme Prepend to the LLM’s input token space (PLITS) integration, a term we have introduced to group many state of the art methods in this family of audio LLMs such as  Wu et al. (2025b); Xu et al. (2025); Chu et al. (2024); Goel et al. (2025); Chu et al. (2023); Ghosh et al. (2024); Tang et al. (2024); Gong et al. (2024); Deshmukh et al. (2023). The second paradigm, Flamingo style architectures (Alayrac et al., 2022; Kong et al., 2024), instead insert cross attention and feedforward (FFN) blocks between successive LLM layers; at each insertion, text tokens attend to a set of latent audio tokens, pass through the block FFN, and only then proceed to the next LLM layer. While this design improves attention efficiency relative to PLITS concatenation, the interleaved cross attention plus FFN stacks increase sequential depth and per layer compute, which can slow the forward pass.
@@
In contrast, we introduce LAL, a lightweight integration that injects audio tokens into the LLM’s attention blocks as keys and values only (without forming audio queries) and bypasses the LLM FFNs for audio tokens. This reduces the attention complexity from $\mathcal{O}\!\big((N_{a}{+}N_{t})^{2}\big)$ to $\mathcal{O}\!\big((N_{a}{+}N_{t})N_{t}\big)$, where $N_{a}$ and $N_{t}$ denote the numbers of audio and text tokens, respectively. Since typically $N_{a}\gg N_{t}$, this yields substantial efficiency gains. By avoiding both quadratic attention over audio tokens and their passage through LLM FFNs, LAL substantially reduces memory usage and computation. Unlike parameter-efficient methods such as LoRA, this is a core architectural modification, so the efficiency benefits are realized not only during training but also at inference time.
@@
Figure 1: Comparison of compute efficiency between PLITS, state of the art audio-LLM integration (our baseline),  Flamingo,  LAL(ours) and  PAL(Ours). Training was performed with batch size 8 on an NVIDIA A100 using bfloat16, and inference with batch size 12 on an NVIDIA A100 using float16. All benchmarks were executed sequentially on the same node to eliminate load-related discrepancies.
LAL provides a compute- and memory-efficient mechanism by constraining how audio tokens interact with the LLM, while achieving performance comparable to PLITS. We further introduce a hybrid variant, PAL, which combines LAL and PLITS to balance efficiency and performance. PAL outperforms PLITS while reducing computational and memory requirements.
@@
To validate these architectural choices, we conduct a systematic empirical study under a standardized training curriculum and dataset setup, ensuring fair comparisons across models. Our experiments explore the trade-off between performance and efficiency, highlighting how different integration techniques facilitate effective information transfer from audio encoders to LLMs with minimal parameter overhead. This analysis provides actionable insights into the design of scalable and efficient audio LLMs.
@@
Our main contributions are as follows: (1) We introduce LAL, a lightweight integration strategy for audio-LLMs that incorporates audio tokens solely as keys and values in the LLM’s attention sub-modules and skips FFNs, thereby reducing computation and memory cost while retaining performance comparable to PLITS integration, (2) We propose PAL, a hybrid integration that applies PLITS to a compact set of summary tokens and LAL to the full audio token sequence, significantly outperforming PLITS while remaining efficient, and (3) We conduct fair and rigorous architectural comparisons under a standardized training curriculum and dataset setup, providing actionable insights into the efficiency–performance trade-offs of audio-LLM design.
@@ 2 Literature review
Audio LLM architectures: When integrating audio encoders with an LLM, two paradigms dominate. In PLITS, encoder features are mapped to the LLM token space with a small projector such as an MLP or a Q Former, the resulting audio tokens are typically prepended to the text tokens, and the joint sequence is processed by all LLM layers (Wu et al., 2025b; Xu et al., 2025; Chu et al., 2024; Goel et al., 2025; Chu et al., 2023; Ghosh et al., 2024; Tang et al., 2024; Gong et al., 2024; Deshmukh et al., 2023). In contrast, the Flamingo style architecture inserts cross attention and feed forward adapters between successive LLM layers so that text tokens attend to latent audio tokens at selected depths (Alayrac et al., 2022; Kong et al., 2024). This makes audio to text interaction explicit and gated, but adds sequential depth, per layer compute, and parameters. An extended literature review is provided in Appendix I.
Figure 2: Illustration of integration techniques: (A) SOTA integration PLITS (prepend to the LLM’s input token space), which prepends audio tokens to text tokens and propagates the full sequence through all LLM layers (our baseline); (B)  Flamingo integration, where text tokens first attend to audio tokens through a separate cross attention plus FFN module, and the resulting signal is added to the text residual stream before the next LLM layer. (C) our proposed lightweight integration LAL, which introduces audio representations only through the attention mechanism (see Equations 2, 3, and 4) while bypassing the feedforward modules; (D) PAL, a hybrid integration that combines LAL and PLITS integrations.
@@ 3 Methodology
This section outlines our approach to integrating audio with language models. We begin by formalizing PLITS, the SOTA audio-LLM integration, as our reference baseline. We then introduce LAL, a lightweight alternative that injects audio through attention only, and we analyze its compute and memory profile. Finally, we connect these findings to PAL, a hybrid integration that applies PLITS to a compact set of summary tokens and LAL to the full audio token sequence, significantly outperforming PLITS while remaining efficient.
@@ 3.1 Baseline Audio LLM: Prepend to the LLM’s input token space (PLITS)
To provide a fair comparison point for our integration methods, we construct a baseline audio LLM that follows the widely adopted SOTA integration strategy, which we refer to as Prepend to the LLM’s input token space (PLITS). In this design, the audio encoder outputs are first mapped into the LLM input embedding space using a Q-Former–style connector. The resulting audio tokens are then prepended to the text tokens, and the concatenated sequence is passed through all LLM layers so that decoding proceeds jointly over audio and text (see Fig. 2(A)).
The central characteristic of this PLITS-style integration is that the audio tokens are prepended to the text tokens. This integration strategy is used by most audio LLMs, including several state of the art systems Wu et al. (2025b); Xu et al. (2025); Chu et al. (2024); Goel et al. (2025); Chu et al. (2023); Ghosh et al. (2024); Tang et al. (2024); Gong et al. (2024); Deshmukh et al. (2023).
@@ 3.2 LAL: Lightweight Audio-LLM Integration
A standard LLM layer consists of an attention submodule followed by a feed-forward network (FFN) submodule. Since attention mediates all inter-token interactions, it is the necessary pathway for audio to influence text, and we posit that it is also sufficient for text tokens to gather information from audio. Guided by this principle, we introduce LAL (Lightweight Audio LLM integration).
As in our baseline, a shared Q-Former produces a sequence of audio tokens and at each layer a small MLP projects these tokens into that layer’s input space. Audio information is then injected into the attention block only through Keys and Values while Queries remain text only, so audio modulates the attention context of text tokens without passing through the feed-forward network.
@@
after which $\tilde{H}_{l}^{t}$ proceeds through the FFN with the usual residual connections. In this way, audio cues shape the attention context seen by text tokens, aligning audio-evoked features with their linguistic counterparts and enabling effective cross–modal information transfer. Please note that in LAL we assign position IDs to both audio and text tokens to preserve token-order information, as detailed in Appendix C and Figure 6.
Information Injection Dynamics. LAL is neither Lite PLITS nor Lite Flamingo. Beyond computational efficiency, LAL opens up a distinct information pathway. In PLITS, audio tokens are treated identically to text tokens: they are transformed layer by layer through causal self-attention and FFN non–linearities, causing their representations to drift from the original encoder output as they mix with the LLM’s internal state. In contrast, LAL uses a dedicated MLP at each layer to project “semantic-ready” audio features directly into the appropriate abstraction for that layer. This preserves a direct link to the audio encoder’s semantic output. For tasks that rely on explicit acoustic cues, such as sound event understanding (e.g., Which animal sound is heard?), this projection-based injection can be more effective than the deeply transformed representations produced by PLITS.
When comparing to Flamingo, we note that although Flamingo also injects semantic level information without decoding audio tokens inside the LLM, the route by which this information influences text tokens is fundamentally different. In Flamingo, text tokens first attend to audio tokens in a dedicated cross attention module; the resulting signal is added to the text residual stream, and only then do the updated text states interact through the standard self attention layers of the LLM. In LAL, by contrast, audio representations are introduced directly into the same self attention operation as the text tokens, so text attends jointly to audio and text within a single attention computation. This produces a distinct information flow from Flamingo. We also note that LAL does not require the extra cross attention plus FFN adapter blocks used in Flamingo.
To summarize, LAL is similar to PLITS in that it performs in-context injection and allows text tokens to attend over both audio and text tokens, and it is similar to Flamingo in that it injects information that has not been fully decoded inside the LLM. However, it is architecturally distinct from both: LAL is neither a “lite PLITS” nor a “lite Flamingo,” but rather a new information pathway for integrating audio encoders with LLMs.
LAL Integration with Frozen LLM FFN. We also verify that LAL integration remains effective when the LLM’s FFN blocks are frozen, with no significant loss in performance (refer to Appendix F). This finding has important implications for reducing training cost, improving parameter efficiency, and preserving the pretrained knowledge of the LLM while enabling multimodal alignment. For clarity and consistency, however, our main experiments focus on the standard setting with trainable FFN blocks, and discussion of the frozen-FFN variant is limited to Appendix F.
Leveraging parametric versus contextual knowledge. Here we posit how LAL efficiently utilizes two types of knowledge inherent in pre-trained LLMs: (1) parametric knowledge, primarily embedded within the FFN layers as a result of extensive language pre-training, and (2) contextual knowledge, which is dynamically incorporated through attention mechanisms. We posit that audio as contextual information can effectively induce required concept activations in text token representations via attention-based modulation, without needing direct FFN processing of audio representations. Consequently, audio information indirectly accesses the LLM’s parametric knowledge: the audio context ”piggybacks” on text tokens, as attention mechanisms reconfigure these representations, which then engage relevant concept-related pathways during FFN processing.
3.2.1 Compute and memory efficiency.
LAL is more compute- and memory-efficient than PLITS and Flamingo style integration, and the benefits become more pronounced with longer audio sequences. At a high level, the gains come from reducing the effective attention complexity and avoiding unnecessary routing of audio tokens through the feed-forward sublayers. Quantitative comparisons of memory usage and training throughput are reported in Figure 1.
In the following subsections, we present one-to-one comparisons between LAL and the PLITS baseline when applicable, and otherwise discuss properties specific to LAL.
Attention Complexity:
@@
LAL: audio tokens do not enter the feedforward sublayer and only serve as keys and values for text queries, which reduces per layer floating point operations and activations stored for backpropagation.
@@
Scaling With Audio Length: Non text modalities in multimodal LLMs often yield far more tokens, and audio is no exception. As $N_{a}$ grows due to longer clips or denser tokenization, PLITS incurs a cost of $(N_{a}+N_{t})^{2}$, so the $N_{a}^{2}$ term dominates. In contrast, LAL scales as $(N_{a}+N_{t})N_{t}$, which is linear in $N_{a}$. Thus, the compute and memory gap widens with longer or more finely segmented audio. The feedforward savings in LAL also increase with $N_{a}$ as a larger share of tokens bypass the most expensive part of each block.
@@
Distinct from PEFT and LoRA: LAL is a core architectural modification, not a parameter-efficient fine-tuning (PEFT) method such as LoRA (Hu et al., 2022). Techniques such as LoRA adjust how weights are adapted during training while keeping the forward compute pattern essentially the same at inference. In contrast, LAL changes how audio tokens participate in attention and feedforward routing, so its compute and memory savings apply at inference as well as during training.
@@
Table 1: Performance evaluation of the proposed efficient integration method LAL and SOTA integration PLITS across different base LLMs. Evaluation follows the protocol of Gong et al. (2024). FI: Flamingo Integration, LFST: a Q-former-based approach that compresses features from SSLAM (fine-grained audio encoder) and CLAP (language-aligned audio encoder) into a single set of Q-former tokens (see Appendix J), AC: AudioCaps, CL: Clotho, AS2M: AudioSet 2M. indicates CIDEr and indicates SPICE. Other metrics: accuracy (ESC-50, VocalSound), Mi-F1 (DCASE), and mAP (FSD, AudioSet). For evaluation methodology see Section 4.1 and for dataset details see Appendix E.
LLM PLITS FI LAL LFST Classification Captioning
@@
Table 2: GPT-4 evaluation of LAL and PLITS on the CompA-R benchmark (Ghosh et al., 2024). LFST: a Q-former-based approach that compresses features from SSLAM (fine-grained audio encoder) and CLAP (language-aligned audio encoder) into a single set of Q-former tokens (see Appendix J). A text only GPT-4 judge scores the model outputs; see Ghosh et al. (2024) for the detailed prompt.